A company that sells different models of socks, faces an annual demand of 10,000 units of its main model. The cost of issuing a purchase order is $ 50 and it has been estimated that the cost of unit storage of the product for one year is $ 2.5. Assume that the provider takes a long time to fill a 7 days order, and the company wants a 5 day security inventory. With a 360-day year, determine:


a) The economic order quantity using EOQ that minimizes total costs.
b) What is the point of reorder?
c) Interpret the resultsType or paste question here
Answer #1
  1. The economic order quantity using EOQ that minimizes total costs.

We have following information

Annual demand D = 10,000 units per year

The cost of issuing a purchase order or Ordering cost S = $50 per order

Annual storage or holding cost H = $2.50 per unit per year

For optimum order quantity per order to minimize the total cost, we have to calculate economic order quantity (EOQ) in following manner

EOQ = sqrt (2* D*S/H)

= sqrt (2*10,000*$50/$2.50) = 632.46 units

Therefore economic order quantity (EOQ) is 632.46 units

  1. What is the point of reorder?

Formula for the Reorder Point calculation

Reorder Point (ROP) = average daily demand * lead time + safety stock

Where,

Reorder Point (ROP) =?

Average daily demand = annual demand / number of days in a year

= $10,000/360 = 27.78 units

Lead time = 7 days

Safety stock = 5 day security inventory = 5* 27.78 units = 138.89 units

Therefore,

Reorder Point (ROP) = 27.78 *7 +138.89 units

= 333.33 units

Reorder Point is 333.33 units

c) Reorder point is the point of inventory or stock in hand when a company placed its next order. The order quantity of the next order will be equal to its economic order quantity (EOQ) so that it can minimize its total cost of ordering and storage.